OpenStudy (anonymous):

The radian measure of an angle that is 125°25'50" is ______. A.) 2.19 radians B.) 2.25 radians C.) 2.20 radians D.) 3.49 radians Please provide an explanation if possible. Thanks. =)

OpenStudy (jdoe0001):

radians = \(\bf \cfrac{degrees \times \pi}{180}\)

OpenStudy (anonymous):

How do I convert 125°25'50" entirely into degrees?

OpenStudy (jdoe0001):

keeping in mind that there are 60 seconds in 1 minute so how many minutes in 50 seconds? well, 50/60

OpenStudy (jdoe0001):

think about it this way, if you had 60 secs, how many minutes are there in 60 seconds? 60/60 = 1min

OpenStudy (jdoe0001):

so, you end up with 50/60 = 0.833 so you augment that to the "25" existing minutes already and you'd end up with 25.833 minutes

OpenStudy (jdoe0001):

so how many degrees in 25.833 minutes? well, there are 60 minutes in 1 degree so you'd have 25.8333/60 degrees you'd end up with another decimal amount, augment that to your degrees

OpenStudy (anonymous):

So 125°25'50" = approx. 125.43°?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

plug that in, and you'd get your radians

OpenStudy (anonymous):

Ah, I get 2.189 so it's A.

OpenStudy (anonymous):

Thanks for your time! =)

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

yw